OpenStudy (anonymous):

(6x)^6 -------- (6x)^18 divide

OpenStudy (anonymous):

1/ (6x)^3

OpenStudy (phi):

The rule is when you have the same base, you subtract exponents. Here your base is 6x so \[\frac{(6x)^6}{(6x)^{18}}= \frac{1}{(6x)^{12}}= (6x)^{-12}\]